OpenStudy (anonymous):

How many kg of pure salt must be added to 8 kg of a 30% solution to obtain a 35% soltution? Round to the nearest tenth.

OpenStudy (radar):

.30 Solution = 8 kg salt solution = 8/.3=26.66kg of salt solution. So we have 8 kg salt and 18.66 kg of something else (probably water). That is what we now have. How much kg of pure salt would be added to produce a 35% solution. Lets call the amount to be added x.

OpenStudy (radar):

8 + x = .35(26.66 +x) 8+x =9.33 +.35x x-.35x=9.33-8 .65x =1.333 x=2.05 kg of salt This should be checked out. By adding 2.05 kg of salt the total salt is now 8+2.05=10.05 kg salt and the total solution is now 26.666 kg + 2.05 kg or a total of 28.7166, but is it a 35% solution.........lets see (.35)(28.7166)=10.05 Hey it checks